OpenStudy (anonymous):

show that integral (ln (sin(x))=-piln(2) from 0..pi.

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi} \ln \sin(x)=-\pi \ln(2)\]

OpenStudy (lgbasallote):

interesting

OpenStudy (anonymous):

ahh this classic integral again...!

OpenStudy (anonymous):

i have solved Numerically . but want to know is there analytical way .?

OpenStudy (anonymous):

yeah there is...

OpenStudy (anonymous):

how ? can you guide me ?

hartnn (hartnn):

i remember solving it....will take some time to type the reply....

hartnn (hartnn):

ok,first since sin is symmetric in the range [0,pi],we can write that as 2*integral from 0 to pi/2,right?.....because there should be mod sign inside ln,as ln cannot take negative values.

hartnn (hartnn):

now we can use the property that \[\int\limits_{a}^{b}f(x)dx=\int\limits_{a}^{b}f(a+b-x)dx\] to get \[\int\limits_{0}^{\pi/2}\ln|\cos x| dx\] with me till here?

OpenStudy (anonymous):

yes. how did you get cos(x)?

hartnn (hartnn):

sin(0+pi/2-x)=cos x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now ?

hartnn (hartnn):

now let the original integral be \[I=2\int\limits_{0}^{\pi/2}\ln|\sin x|=2\int\limits_{0}^{\pi/2}\ln|\cos x|\] so can i write that as \[I=\int\limits_{0}^{\pi/2}\ln|sinx|dx+\int\limits_{0}^{\pi/2}\ln|\cos x|dx\] be careful,this is the main step

OpenStudy (anonymous):

ok .

hartnn (hartnn):

so now,i would use definite integral and ln properties,like this \[I=\int\limits_{0}^{\pi/2}\ln|\sin x \cos x|=\int\limits_{0}^{\pi/2}\ln|2\sin x \cos x|-\ln 2\] got this? then u can write 2 sin x cos x as sin 2x and make the substitution of 2x=t,try it and tell me what u get..

hartnn (hartnn):

sorry,i keep on forgetting dx.....

OpenStudy (anonymous):

that's a beautiful proof.

OpenStudy (anonymous):

i am getting \[\int\limits_{0}^{\pi/2}\ln|\sin(2x)-\ln(2)\] \[\int\limits_{0}^{\pi}[\frac{1}{2}\ln|\sin(t)-ln(2)]dt\]

hartnn (hartnn):

absolutely correct,now separate those two integrals,solve 2nd one(ln 2) and in 1st integral,write \[I=\int\limits_{0}^{\pi}\ln|\sin x|\] and don't forget to write ,I=.............. then u can solve for I to get to your answer tell me if this was confusing....

OpenStudy (anonymous):

should i solve them separately ?

hartnn (hartnn):

the 1st integral need not be solved,u get 1/2 I ,and 2nd integral when solved will just give pi/2*ln2 ,right? so now u have I=I/2-pi/2*ln 2 solve for I

OpenStudy (anonymous):

ohh thanks .. i got it now.

hartnn (hartnn):

welcome :) its good to remember such beautiful proofs......

OpenStudy (anonymous):

there is a lot of beautiful integrals in this putnam competitions http://mks.mff.cuni.cz/kalva/putnam/putn53.html

OpenStudy (anonymous):

http://mks.mff.cuni.cz/kalva/putnam.html

hartnn (hartnn):

and u have solved them all??

OpenStudy (anonymous):

lol...most of them